Which is an equation for the parabola with focus at (-11,-4) and vertex at (1,-4)
Answer by lwsshak3(11628)

Which is an equation for the parabola with focus at (-11,-4) and vertex at (1,-4)
**
This is an equation of a parabola that opens leftwards
Its standard form: (y-k)^2=-4p(x-h), (h,k)=(x,y) coordinates of vertex.
axis of symmetry: y=-4
p=12 (distance from vertex to focus on he axis of symmetry. ( from 1 to -11)
4p=48
equation:
(y+4)^2=-48(x-1)